From mar@roe.ac.uk Fri Jul 11 16:08:52 2014 Date: Fri, 11 Jul 2014 15:16:27 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Dave Morris , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 04 July 2014 WFAU VDFS Science Archives progress meeting minutes, 04 July 2014 Present: MAR, RPB, ETWS, RSC, ACD Apologies: NJC, RGM, NCH, KTN, AL, MSH DoNM: 10am Friday 18 July 2014 in the Villas meeting room, ACD in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: ETWS to investigate modifying post CU4 ops to calculate and store illumination correction values. Done and tested. ACTION: RPB to create a wiki page to discuss ways of generating unique objIDs. Done and the options would be discussed at the next meeting. ACTION: RPB/ACD/MSH to see if a new database node could be ordered. Done, approval given and machine had been ordered. ACTION: MSH, RPB conduct a tape inventory to check if any tapes are missing. Done, none are missing. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 20/6/2014 meeting: ---------------------------------------------- Action: CTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Continues. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan). ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: MAR to document issue with ATLAS unpaired sources and give examples of recovering matches. Continues. ACTION: ACD to arrange follow-up VVV brainstorming meeting. Continues. RSC was organising the testing and the meeting should be organised ICW him. ACTION: RPB to write a script to backup the SVN repository. Continues ACTION: RPB estimate the number of additional tapes that would be needed to backup all the data on disk. Continues ACTION: NJC to circulate the relevant VPHAS related e-mail messages. Continues, NJC not present ACTION NJC: update ESO VISTA releases on the deadlines web page. Continues, NJC not present ACTION: MSH to produce a costed list of the additional kit necessary to permit wholesale reprocessing of UKIDSS/VISTA. Continues. ACTION: ACD, RGM: review finances. Continues Specific points and new actions: ------------------------------- Project management: Nothing to report. WFCAM, VISTA and VST updates: The telescopes were operating normally. Comments and issues arising from CASU minutes: No minutes had been received since the previous meeting. Review of Deadlines: The GES DR2 release was well behind schedule but was now available to beta testers. ETWS noted that database space was needed to progress OSA ingests and an ATLAS APASS release. ACTION: RPB to sort out disk space for the new OSA ingest database. Networking: ETWS noted that speeds had been quite variable recently but the average is good enough for us keep up-to-date with transfers. WSA/VSA/OSA/GES Operations: ETWS reported that VSA CU1-CU2 was complete up to April 2014. OSA v1.1 transfers were complete and would be ingested into a new database. January-March 2014 WFCAM data had been made available by Mike Irwin and were being transferred. RSC asked about the GPS DR10 release. MAR admitted he had failed to get back to closing out the QC. ACTION: MAR to finish off GPS QC and add entry to forthcoming deadlines. RPB reported that the Spitzer SMC catalogue requested by the VMC team had been downloaded and added to the Spitzer database. ACTION: RPB to make the necessary Spitzer SMC/VMC neighbour table entries. Hardware and Systems: The new database node had been ordered. The Nytro machine had gone back again. MSH was in the process of getting quotes for and comparing 10Gb ethernet and upgraded infiniband. ETWS noted that the latest NAS box was available as disks 47-51. Software: ACD and RSC had been working on tweaks to the GES ingest/release. ETWS reported he'd been working on the OSA, ingesting corrected, uncorrected and list driven catalogues. This had been tested successfully on a night's worth of data. Survey Data Release: VIKING and VMC ESO release files had been prepared and were in the process of being uploaded. A GES iDR2 beta release had been made and feedback was coming in. Non-survey Data Release: No releases had been made. Astrogrid deployment, VO & Data Analysis services: Nothing for the minutes. Miscellaneous: ADASS submissions are being sought. -- Scanned by iCritical.